for reaction N 2 + 3H 2 gives 2NH 3 + heat a] Kp = Kc b] Kp = Kc * RT c] Kc = Kp * (RT) -2 d] Kc = Kp (RT) -1

for reaction N2 + 3H2  gives 2NH3    + heat


a]    Kp =  Kc              b]     Kp = Kc *  RT


c]     Kc =  Kp * (RT) -2        d]   Kc = Kp (RT)  -1                       

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear dhawal.

Kp=KcRTn

Here m=2-4=-2

therefore the correct answer should be Kp=KcRT-2

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi

vikas askiitian expert
13 years ago

kp=kc.(RT)^-2